OpenStudy (anonymous):

STATISTICS HELP!!! WILL FAN!!! Sara is reviewing recent orders at her deli to determine which meats she should order. She found that of 1,000 orders, 390 customers ordered turkey, 345 customers ordered ham, and 300 customers ordered neither turkey nor ham. Based on this data, how many of the next 1,000 customers will order both turkey and ham? Show your work, and use complete sentences. I don't get why when you add up those number they add to a number above 1,000..

OpenStudy (anonymous):

@xapproachesinfinity @micheleee13 @isntshelovelyy

OpenStudy (anonymous):

@NinjaDevo @Norbyboo

OpenStudy (anonymous):

is it multiple choise..?

OpenStudy (anonymous):

No it's written :/

OpenStudy (anonymous):

I'm confused why they don't add up to 1000 too... I'm not in statistics either...maybe someone with more knowledge can help. :)

OpenStudy (anonymous):

I'm actually in algebra 2 but the chapter is on statistics :/ @NinjaDevo

OpenStudy (anonymous):

i just re posted this if i get the awnser ill let u knowso dont delete this

OpenStudy (anonymous):

@isntshelovelyy thank you!!

OpenStudy (anonymous):

yep :) np

OpenStudy (anonymous):

@sierraurb Someone answered it here: http://openstudy.com/study#/updates/53ec28dae4b01789aba4d1fc

OpenStudy (anonymous):

@NinjaDevo @isntshelovelyy shouldn't there be a formula or something?

OpenStudy (anonymous):

FLVS huh?

OpenStudy (anonymous):

@vineeth10 yup lol

OpenStudy (anonymous):

i did that already...let me see if i can find it

OpenStudy (anonymous):

what module is it?

OpenStudy (anonymous):

@vineeth10 9!! in the written part of the test :)

OpenStudy (anonymous):

i JUST finished it lol

OpenStudy (anonymous):

Module test?

OpenStudy (anonymous):

@vineeth10 yes!! thank you soooo much

OpenStudy (anonymous):

550+305+225=1080. 1080-1000=80. So, 80 people will order both ham and turkey. Yours is just different numbers

OpenStudy (anonymous):

@vineeth10 thank you so so so so SO much!! you're seriously a lifesaver!! :)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

just be careful cause mine is different numbers

OpenStudy (anonymous):

@vineeth10 I will!! i was mainly confused as to why it added up to be over 1000

OpenStudy (anonymous):

oh ok lol

OpenStudy (science0229):

Sorry to interrupt, but you can also use set theory for this I put the amount of orders for both ham and turkey as x. All of them combined should be 1000. 300+390-x+x+345-x=1000 1035-x=1000 Therefore, 35 customers will order both ham and turkey.

OpenStudy (anonymous):

U are right @science0229 that is the answer for @sierraurb

OpenStudy (anonymous):

@science0229 thank you so much!! that's what i put for my answer :)

OpenStudy (anonymous):

i have flvs to i hatee it lol but somebody awnserd it on mine